Consider a firm facing a linear demand function and a constant marginal cost given below:

𝑝(𝑄)=π‘Žβˆ’π‘π‘„ and 𝑐(𝑄)=𝑐𝑄 where c is the constant marginal cost.

Find out the profit maximizing level output produced by the firm. Also, check the second order necessary condition for the same.


1
Expert's answer
2021-03-08T09:22:03-0500

a) The profit maximizing level output produced by the firm occurs when MR=MC. Let's first find the total revenue:


"TR=PQ=(a-bQ)Q=aQ-bQ^2."

Then, we can find the marginal revenue:


"MR=\dfrac{dTR}{dQ}=a-2bQ."

By the definition of the marginal cost, we have:


"MC=\dfrac{dTC}{dQ}=c."

Finally, we can find the profit maximizing level output produced by the firm:


"a-2bQ=c,""Q=\dfrac{a-c}{2b}."

b) The second order condition can be witten as follows:


"\dfrac{d^2TR}{dQ}-\dfrac{d^2TC}{dQ}<0,""\dfrac{d}{dQ}(a-2b)-\dfrac{d}{dQ}(c)<0,""-2b<0."

Therefore, as we can see, the second order condition is satisfied.